A Carnot engine is operated between two heat reservoirs at T1= 435 K and T2= 308 K. (a) If the engine receives 5113 J of heat from the reservoir at T1 in each cycle, how many joules per cycle does it deliver to the reservoir at T2? (b) How much mechanical work is performed by the engine during each cycle? (c) What is the thermal efficiency of the engine?
Answer #1

c) thermal efficiency = 1 - Tc/Th = 1 - (308/435) = 0.2919 = 29.19%

b) now efficiency also equals = W/Qh = 0.2919

=> W = 0.2919*5113 = 1492.48 J

a) heat delivered to cold reservoir

Qc = Qh - W = 5113 - 1492.48 = 3620.51 J